OpenStudy (anonymous):

Verify the following identity: 1-2cos^2(x) = 2sin^2(x) - 1 Please include steps! :/ Thank you!

OpenStudy (anonymous):

Sin^2 x + cos^2 x = 1

OpenStudy (lalaly):

1-2(1-sin^2x) = 1-2+2sin^2x = 2sin^2x-1 :D

OpenStudy (anonymous):

Thank you!!